# Copresheaf Topological Neural Networks: A Generalized Deep Learning Framework

### v2

@@ B.4 Copresheaf Neighborhood Matrix Example
Appendix C Expressive Power of CTNNs
C.1 Universal Approximation of $\mathcal{N}$-Dependent Copresheaves
@@
Proposition 5 (Universal approximation of $\mathcal{N}$-dependent copresheaves).
@@
be given so that the $2d$-vectors $(\mathbf{h}_{y},\mathbf{h}_{x})$ are pairwise distinct for every directed edge $y\to x$. Define
@@
Since $A$ is finite and its elements are distinct, the assignment $g\colon A\to\mathbb{R}^{d\times d}$ is well-defined.
@@
Appendix D Sheaf Neural Networks Are Copresheaf Message-Passing Neural Networks
In this appendix, we prove Theorem 1 and Proposition 2, which demonstrate that existing sheaf neural networks (SNNs), including sheaf diffusion networks, are special cases of the copresheaf message-passing neural network (CMPNN) framework (Definition 1). Moreover, we summarize how other sheaf-based neural architectures align with our unifying message-passing framework (see Table LABEL:tab:sheaf-homp-summary).
@@ Appendix D Sheaf Neural Networks Are Copresheaf Message-Passing Neural Networks
First, construct the bidirected graph $G^{\prime}=(V,E^{\prime})$ from $G=(V,E)$ by replacing each undirected edge $\{u,v\}\in E$ with two directed edges $(u,v),(v,u)\in E^{\prime}$. This ensures that $G^{\prime}$ retains the connectivity of $G$ while introducing explicit directionality.
@@
  • For each vertex $x\in V$, assign $\mathcal{G}(x)=\mathcal{F}(x)$, where $\mathcal{F}(x)$ is the vector space associated to $x$ by the cellular sheaf $\mathcal{F}$.
  • For each directed edge $(v,u)\in E^{\prime}$, corresponding to the undirected edge $e=\{u,v\}\in E$, define the linear morphism $\rho_{v\to u}:\mathcal{G}(v)\to\mathcal{G}(u)$ by $\rho_{v\to u}=\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}$, where $\mathcal{F}_{v\triangleleft e}:\mathcal{F}(v)\to\mathcal{F}(e)$ and $\mathcal{F}_{u\triangleleft e}:\mathcal{F}(u)\to\mathcal{F}(e)$ are the restriction maps of $\mathcal{F}$, and $\mathcal{F}_{u\triangleleft e}^{\top}:\mathcal{F}(e)\to\mathcal{F}(u)$ is the adjoint with respect to an inner product on $\mathcal{F}(e)$.
Now, consider the SNN message-passing mechanism along the edge $e=\{u,v\}$. For a feature vector $\mathbf{h}_{v}\in\mathcal{F}(v)$ at vertex $v$, the message transmitted to vertex $u$ is given by $\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}(\mathbf{h}_{v})$.
In the copresheaf $\mathcal{G}$ on $G^{\prime}$, the morphism associated with the directed edge $(v,u)$ is $\rho_{v\to u}=\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}$. Applying this morphism, the message-passing operation in the CMPNN yields $\rho_{v\to u}(\mathbf{h}_{v})=\mathcal{F}_{u\triangleleft e}^{\top}\circ\mathcal{F}_{v\triangleleft e}(\mathbf{h}_{v})$, which is identical to the SNN message.
To ensure that $\mathcal{G}$ is a well-defined copresheaf, observe that it assigns vector spaces to vertices and linear maps to directed edges in a functorial manner. Specifically, for each directed edge $(v,u)\in E^{\prime}$, the map $\rho_{v\to u}$ is a composition of linear maps and thus linear. The identity and composition properties are satisfied implicitly through the consistency of the sheaf restriction maps.
@@
$(I_{n}\otimes W_{1})\mathbf{H}=[W_{1}\mathbf{h}_{v}]_{v\in V},$
$\bigl{(}\Delta_{\mathcal{F}}\otimes I\bigr{)}(I_{n}\otimes W_{1})\mathbf{H}W_{2}=\left[\sum_{u\in V}L_{F,v,u}W_{1}\mathbf{h}_{u}W_{2}\right]_{v\in V},$
$\mathbf{h}_{v}^{+}=\mathbf{h}_{v}-\sum_{u\in\mathcal{N}(v)\cup\{v\}}W_{2}L_{F,v,u}W_{1}\mathbf{h}_{u},$
since $L_{F,v,u}=0$ for $u\notin\mathcal{N}(v)\cup\{v\}$.
Interpret $G$ as a directed graph with edges $u\to v$ for $u\in\mathcal{N}(v)$ and $v\to v$. Define: - Message function: $\alpha(\mathbf{h}_{v},\rho_{u\to v}\mathbf{h}_{u})=W_{2}L_{F,v,u}W_{1}\mathbf{h}_{u}$, - Morphisms: $\rho_{u\to v}$ implicitly encoded via $L_{F,v,u}$, - Aggregator: $\oplus$=$\sum$ = ∑, - Update: $\beta(\mathbf{h}_{v},m)=\mathbf{h}_{v}-m$.
@@
$\mathbf{h}_{v}^{+}=\beta\left(\mathbf{h}_{v},\sum_{u\to v}\alpha(\mathbf{h}_{v},\rho_{u\to v}\mathbf{h}_{u})\right),$
@@
Table LABEL:tab:sheaf-homp-summary provides a summary of sheaf neural networks realized in terms of Definition 1.

### v3

@@ B.4 Copresheaf Neighborhood Matrix Example
Appendix C Copresheaf Laplacian, Energy, and CTNN Transport–diffusion
In this section we introduce a linear transport-discrepancy operator $B_{\rho}$ that maps node fields to edge discrepancies, define the quadratic energy $E_{\rho}$ and the associated Laplacian $L_{\rho}=B_{\rho}^{\top}WB_{\rho}$, derive its block form and kernel, and show that a CTNN residual layer implements an explicit gradient step that monotonically decreases $E_{\rho}$ for suitable step sizes. This energy perspective provides an interpretable and theoretical foundation for the CTNN architecture, framing it as a diffusion process that minimizes a specific energy function. Our copresheaf Laplacian also coincides with the quiver Laplacian [Sumray et al., 2024] when viewing the transport maps as arrow representations, connecting our framework to established spectral methods while enabling directional, anisotropic information flow distinct from the symmetric diffusion in sheaf-based models Bodnar et al. [2022].
Spaces and operator
Let $(F,\rho,G)$ be a copresheaf defined on a directed graph $G=(V,E)$. To each node $x\in V$ attach a finite-dimensional real inner-product space $F(x)$. For every edge $y\to x\in E$ fix a linear transport $\rho_{y\to x}:F(y)\to F(x)$. Define the node and edge product spaces
$\mathcal{H}_{V}\;\coloneqq\;\bigoplus_{x\in V}F(x),\qquad\mathcal{H}_{E}\;\coloneqq\;\bigoplus_{(y\to x)\in E}F(x),$
equipped with canonical (blockwise) inner products $\langle\mathbf{h},\tilde{\mathbf{h}}\rangle_{V}=\sum_{x\in V}\langle\mathbf{h}_{x},\tilde{\mathbf{h}}_{x}\rangle$ and $\langle\bm{\xi},\tilde{\bm{\xi}}\rangle_{E}=\sum_{(y\to x)\in E}\langle\bm{\xi}_{y\to x},\tilde{\bm{\xi}}_{y\to x}\rangle.$
Definition 14 (Transport-discrepancy operator).
The linear operator
$B_{\rho}:\ \mathcal{H}_{V}\longrightarrow\mathcal{H}_{E}$
is defined componentwise by
$(B_{\rho}\mathbf{h})_{(y\to x)}\;=\;\mathbf{h}_{x}\;-\;\rho_{y\to x}\,\mathbf{h}_{y}\;\in F(x),\qquad\forall\,(y\to x)\in E.$ (8)
Let $B_{\rho}^{\top}:\mathcal{H}_{E}\to\mathcal{H}_{V}$ denote the Euclidean adjoint. A direct calculation yields
$(B_{\rho}^{\top}\bm{\xi})_{x}\;=\;\sum_{y\to x}\bm{\xi}_{y\to x}\;-\;\sum_{x\to z}\rho_{x\to z}^{\top}\,\bm{\xi}_{x\to z},\qquad\forall\,x\in V,\ \bm{\xi}\in\mathcal{H}_{E}.$ (9)
Definition 15 (Energy and copresheaf Laplacian).
Let $w_{y\to x}>0$ be edge weights and define the diagonal operator $W:\mathcal{H}_{E}\to\mathcal{H}_{E}$ by $(W\bm{\xi})_{y\to x}=w_{y\to x}\,\bm{\xi}_{y\to x}$. The weighted transport energy and the copresheaf Laplacian are
$E_{\rho}(\mathbf{h})\;=\;\|B_{\rho}\mathbf{h}\|_{W}^{2}\;=\;\sum_{(y\to x)\in E}w_{y\to x}\,\|\mathbf{h}_{x}-\rho_{y\to x}\mathbf{h}_{y}\|^{2},\qquad L_{\rho}\;=\;B_{\rho}^{\top}\,W\,B_{\rho}.$ (10)
Remark 2 (Relation to quiver Laplacian).
Interpreting the transports $\rho_{y\to x}$ as arrow maps of a quiver representation, the copresheaf Laplacian $L_{\rho}$ in Equation 10 coincides with the standard (weighted) quiver Laplacian $B^{\top}WB$ of that representation. See Sumray et al. [2024] for more about the quiver Laplacian.
It is sometimes useful to have the copresheaf Laplacian in its exact nodewise form. Namely, for every outgoing edge $x\to z$ with map $\rho_{x\to z}$ and weight $w_{x\to z}$, include the reverse edge $z\to x$ with $\rho_{z\to x}:=\rho_{x\to z}^{\top}$ and $w_{z\to x}:=w_{x\to z}$. With this convention one has the nodewise form of the copresheaf Laplacian:
$(L_{\rho}\mathbf{h})_{x}=\sum_{y\in N(x)}w_{y\to x}\big(\mathbf{h}_{x}-\rho_{y\to x}\mathbf{h}_{y}\big).$ (11)
Theorem 2.
With $w_{y\to x}>0$, $L_{\rho}=B_{\rho}^{\top}WB_{\rho}:\mathcal{H}_{V}\to\mathcal{H}_{V}$ is symmetric positive semidefinite and $\mathbf{h}^{\top}L_{\rho}\mathbf{h}=\|B_{\rho}\mathbf{h}\|_{W}^{2}\geq 0$. Its block action at node $x$ is
$(L_{\rho}\mathbf{h})_{x}=\sum_{y\to x}w_{y\to x}\,\bigl(\mathbf{h}_{x}-\rho_{y\to x}\mathbf{h}_{y}\bigr)\;+\;\sum_{x\to z}w_{x\to z}\,\rho_{x\to z}^{\top}\bigl(\rho_{x\to z}\mathbf{h}_{x}-\mathbf{h}_{z}\bigr).$
Moreover,
$\ker L_{\rho}\;=\;\ker B_{\rho}\;=\;\bigl\{\mathbf{h}\in\mathcal{H}_{V}:\;\mathbf{h}_{x}=\rho_{y\to x}\mathbf{h}_{y}\ \text{for all }y\to x\in E\bigr\}.$
Proof.
P.s.d. and the quadratic identity follow from $L_{\rho}=B_{\rho}^{\top}WB_{\rho}$ with $W\succ 0$. The block formula follows by expanding $B_{\rho}^{\top}(WB_{\rho}\mathbf{h})$ via 10. Finally, $\|B_{\rho}\mathbf{h}\|_{W}^{2}=0\iff B_{\rho}\mathbf{h}=0$. ∎
C.1 CTNN/CMPNN residual as copresheaf diffusion.
From the copresheaf message passing equation in Definition 9, choose the edge message $\alpha(\mathbf{h}_{x},\rho_{y\to x}\mathbf{h}_{y})=\mathbf{h}_{x}-\rho_{y\to x}\mathbf{h}_{y}$, sum aggregation, and residual update $\beta(\mathbf{h}_{x},m)=\mathbf{h}_{x}-\eta m$. Then a CTNN/CMPNN layer updates
$\mathbf{h}_{x}^{(\ell+1)}=\mathbf{h}_{x}^{(\ell)}-\eta\!\!\sum_{y\in N(x)}\!\!w_{y\to x}\,\Bigl(\mathbf{h}_{x}^{(\ell)}-\rho_{y\to x}\mathbf{h}_{y}^{(\ell)}\Bigr)\;=\;\bigl(I-\eta L_{\rho}\bigr)\,\mathbf{h}^{(\ell)}.$ (14)
We have the following theorem.
Theorem 3.
For $E_{\rho}$ in $(13)$,
$\nabla E_{\rho}(\mathbf{h})=2\,L_{\rho}\mathbf{h},\qquad\mathbf{h}^{(\ell+1)}=\mathbf{h}^{(\ell)}-\eta\,\nabla\!\Big(\tfrac{1}{2}E_{\rho}\Big)\!\big(\mathbf{h}^{(\ell)}\big).$
Hence $(14)$ is an explicit gradient step on $\tfrac{1}{2}E_{\rho}$. If $0<\eta<\|L_{\rho}\|_{2}^{-1}$, then $E_{\rho}(\mathbf{h}^{(\ell+1)})\leq E_{\rho}(\mathbf{h}^{(\ell)})$, with strict inequality whenever $\mathbf{h}^{(\ell)}\notin\ker L_{\rho}$. Moreover, $I-\eta L_{\rho}$ is non-expansive in $\|\cdot\|_{2}$ and strictly contractive on $\ker L_{\rho}^{\perp}$, so $\mathbf{h}^{(\ell)}$ converges to the orthogonal projection of $\mathbf{h}^{(0)}$ onto $\ker L_{\rho}$.
Proof.
Since $E_{\rho}(\mathbf{h})=(B_{\rho}\mathbf{h})^{\top}W(B_{\rho}\mathbf{h})$, the chain rule gives $\nabla E_{\rho}(\mathbf{h})=2B_{\rho}^{\top}WB_{\rho}\mathbf{h}=2L_{\rho}\mathbf{h}$, so $(14)$ is gradient descent on $\tfrac{1}{2}E_{\rho}$. Let $L_{\rho}=U\Lambda U^{\top}$ with $\Lambda=\operatorname{diag}(\lambda_{i}\geq 0)$ and write $\mathbf{h}^{(\ell)}=U\mathbf{c}^{(\ell)}$. Then $c_{i}^{(\ell+1)}=(1-\eta\lambda_{i})c_{i}^{(\ell)}$ and $E_{\rho}(\mathbf{h}^{(\ell)})=\sum_{i}\lambda_{i}(c_{i}^{(\ell)})^{2}$. If $0\leq\eta\lambda_{i}\leq 1$ for all $i$, then $\lambda_{i}(1-\eta\lambda_{i})^{2}\leq\lambda_{i}$, giving monotone decay, strict if some $\lambda_{i}>0$ has $c_{i}^{(\ell)}\neq 0$. Non-expansiveness and convergence follow from $|1-\eta\lambda_{i}|\leq 1$ (and <1<1 for $\lambda_{i}>0$). ∎
Remark 3.
$\rho_{y\to x}$ may be (i) direct per-edge linear maps, (ii) factored as $\rho_{y\to x}=F_{x}^{\top}F_{y}$ to reduce parameters, or (iii) constrained (e.g. softly orthogonal via $\|\rho_{y\to x}^{\top}\rho_{y\to x}-I\|_{F}^{2}$) to regularize the spectrum of $L_{\rho}$. A learnable (possibly per-layer/head) stepsize $\eta$, normalized by a running estimate of $\|L_{\rho}\|_{2}^{-1}$, enforces the energy-decay guarantee in Theorem 3. When $\rho_{y\to x}=I$, $(13)$ reduces to the (vector-valued) graph Laplacian; orthogonal $\rho$ recovers a connection Laplacian.
Appendix D Expressive Power of CTNNs
D.1 Universal Approximation of $\mathcal{N}$-Dependent Copresheaves
@@
Proposition 2 (Universal approximation of $\mathcal{N}$‑dependent copresheaves).
@@
be given so that the 2d2d‑dimensional vectors $(\mathbf{h}_{y},\mathbf{h}_{x})$ are pairwise distinct for every directed edge $y\to x$. Define
@@
Since $A$ is finite and its elements are distinct, the assignment $g\colon A\to\mathbb{R}^{d\times d}$ is well‑defined.
@@
Appendix E Sheaf Neural Networks Are Copresheaf Message-Passing Neural Networks
The computational use of a cellular sheaf on a graph rests on the incidence poset. Let $G=(V,E)$ and consider the poset on $V\!\cup\!E$ with $x\preceq e$ whenever $x\in e$. A cellular sheaf $\mathcal{F}$ assigns a vector space to each cell and a linear structure map to each incidence. In our convention, the “restriction” along $x\preceq e$ is implemented as a vertex-to-edge lift $\mathcal{F}_{x\trianglelefteq e}:\mathcal{F}(x)\to\mathcal{F}(e)$. Equipping $\mathcal{F}(e)$ with an inner product yields the adjoint $\mathcal{F}_{x\trianglelefteq e}^{\top}:\mathcal{F}(e)\to\mathcal{F}(x)$, a canonical edge-to-vertex back-projection. Message passing between adjacent vertices then arises by composing incidence maps: for $e=\{x,y\}$, the message from $y$ to $x$ is
$\mathcal{F}_{x\trianglelefteq e}^{\top}\circ\mathcal{F}_{y\trianglelefteq e}:\mathcal{F}(y)\to\mathcal{F}(x).$
This composition induces a direction of information flow on an undirected graph while remaining faithful to the sheaf poset structure. Diffusion-style updates and sheaf-Laplacian operators are recovered by aggregating such edge-mediated messages over $\mathcal{N}(x)$. See Figure 6 for an illustration.
Figure 6: Sheaf-induced message passing on an edge $e=\{x,y\}$. (a) Local spaces $\mathcal{F}(x),\mathcal{F}(y),\mathcal{F}(e)$ with vertex-to-edge lifts $\mathcal{F}_{x\trianglelefteq e}$ and $\mathcal{F}_{y\trianglelefteq e}$ along the incidences $x\preceq e$, $y\preceq e$. (b) Incidence-poset view: arrows encode the sheaf’s linear maps attached to $x\preceq e$ and $y\preceq e$. (c) Edge-mediated message passing: the inner product on $\mathcal{F}(e)$ yields adjoints $\mathcal{F}_{x\trianglelefteq e}^{\top}$. The message from $y$ to $x$ is the composition $\mathcal{F}_{x\trianglelefteq e}^{\top}\circ\mathcal{F}_{y\trianglelefteq e}$, i.e., a direct vertex-to-vertex morphism $\rho_{y\to x}:\mathcal{F}(y)\to\mathcal{F}(x)$, compatible with the bidirected expansion $G^{\prime}$. This realizes SNN message passing as CMPNN morphisms assembled from sheaf structure maps.
In this appendix, we prove Theorem 1 and Proposition 1, which demonstrate that existing sheaf neural networks (SNNs), including sheaf diffusion networks, are special cases of the copresheaf message-passing neural network (CMPNN) framework (Definition 9). Moreover, we summarize how other sheaf-based neural architectures align with our unifying message-passing framework (see Table 8).
@@ Appendix E Sheaf Neural Networks Are Copresheaf Message-Passing Neural Networks
First, construct the bidirected graph $G^{\prime}=(V,E^{\prime})$ from $G=(V,E)$ by replacing each undirected edge $\{x,y\}\in E$ with two directed edges $(x,y),(y,x)\in E^{\prime}$. This ensures that $G^{\prime}$ retains the connectivity of $G$ while introducing explicit directionality.
@@
  • For each vertex $x\in V$, assign $\mathcal{G}(x)=\mathcal{F}(x)$, where $\mathcal{F}(x)$ is the vector space associated with $x$ by the cellular sheaf $\mathcal{F}$.
  • For each directed edge $(y,x)\in E^{\prime}$, corresponding to the undirected edge $e=\{x,y\}\in E$, define the linear morphism $\rho_{y\to x}:\mathcal{G}(y)\to\mathcal{G}(x)$ by $\rho_{y\to x}=\mathcal{F}_{x\triangleleft e}^{\top}\circ\mathcal{F}_{y\triangleleft e}$, where $\mathcal{F}_{y\triangleleft e}:\mathcal{F}(y)\to\mathcal{F}(e)$ and $\mathcal{F}_{x\triangleleft e}:\mathcal{F}(x)\to\mathcal{F}(e)$ are the restriction maps of $\mathcal{F}$, and $\mathcal{F}_{x\triangleleft e}^{\top}:\mathcal{F}(e)\to\mathcal{F}(x)$ is the adjoint with respect to an inner product on $\mathcal{F}(e)$.
Now, consider the SNN message-passing mechanism along the edge $e=\{x,y\}$. For a feature vector $\mathbf{h}_{y}\in\mathcal{F}(y)$ at vertex $y$, the message transmitted to vertex $x$ is given by $\mathcal{F}_{x\triangleleft e}^{\top}\circ\mathcal{F}_{y\triangleleft e}(\mathbf{h}_{y})$.
In the copresheaf $\mathcal{G}$ on $G^{\prime}$, the morphism associated with the directed edge $(y,x)$ is $\rho_{y\to x}=\mathcal{F}_{x\triangleleft e}^{\top}\circ\mathcal{F}_{y\triangleleft e}$. Applying this morphism, the message-passing operation in the CMPNN yields $\rho_{y\to x}(\mathbf{h}_{y})=\mathcal{F}_{x\triangleleft e}^{\top}\circ\mathcal{F}_{y\triangleleft e}(\mathbf{h}_{y})$, which is identical to the SNN message.
To ensure that $\mathcal{G}$ is a well-defined copresheaf, observe that it assigns vector spaces to vertices and linear maps to directed edges in a functorial manner. Specifically, for each directed edge $(y,x)\in E^{\prime}$, the map $\rho_{y\to x}$ is a composition of linear maps and thus linear. The identity and composition properties are satisfied implicitly through the consistency of the sheaf restriction maps.
@@
$(I_{n}\otimes W_{1})\mathbf{H}$ $=[W_{1}\mathbf{h}_{x}]_{x\in V},$
$\bigl(\Delta_{\mathcal{F}}\otimes I\bigr)(I_{n}\otimes W_{1})\mathbf{H}W_{2}$ $=\left[\sum_{y\in V}L_{F,x,y}\,W_{1}\mathbf{h}_{y}\,W_{2}\right]_{x\in V},$
$\mathbf{h}_{x}^{+}$ $=\mathbf{h}_{x}-\sum_{y\in\mathcal{N}(x)\cup\{x\}}W_{2}\,L_{F,x,y}\,W_{1}\mathbf{h}_{y},$
since $L_{F,x,y}=0$ for $y\notin\mathcal{N}(x)\cup\{x\}$.
Interpret $G$ as a directed graph with edges $y\to x$ for $y\in\mathcal{N}(x)$ and $x\to x$. Define:
  • Message function: $\alpha(\mathbf{h}_{x},\rho_{y\to x}\mathbf{h}_{y})=W_{2}\,L_{F,x,y}\,W_{1}\mathbf{h}_{y}$,
  • Morphisms: $\rho_{y\to x}$ implicitly encoded via $L_{F,x,y}$,
  • Aggregator: $\oplus=\sum$,
  • Update: $\beta(\mathbf{h}_{x},m)=\mathbf{h}_{x}-m$.
@@
$\mathbf{h}_{x}^{+}=\beta\!\left(\mathbf{h}_{x},\sum_{y\to x}\alpha(\mathbf{h}_{x},\rho_{y\to x}\mathbf{h}_{y})\right),$
@@
Table 8 provides a summary of sheaf neural networks realized in terms of Definition 9.